Please show me how to do these hard homework problems:
1. if Pr(E)= 0.59 then Pr(E')=?
2. If Pr(E)= 0.6, Pr(EUF)= 0.8, Pr(E∩F)= 0.25, then Pr(F)=?
1.05= 0.6 + F
F= .99
3. S= {a,b,c,d,e}, Pr{a,b,c}= 0.55, Pr{c}= 0.15, Pr{c,d}= 0.35, Then Pr{e}=?
0.55 + 0.35 - 0.15 =
e= .75
Please show me if I did these correctly

Answer by kingme18(98)

1. E and E' are complementary events. So, if E happens, then E' doesn't, and if E' happens, then E doesn't. Thus, if Pr(E) = 0.59, then Pr(E') = 1 - 0.59 = 0.41
2. You were almost right. The only problem you had is that 1.05 - 0.6 = 0.45, not 0.99. You did 1.05 - 0.06...double check your decimal placement :)
3. You found Pr{a,b,c,d}. Pr{e} is the leftover: 1 - 0.75 = 0.25.